OpenStudy (anonymous):

Find all points of discontinuity of the given function. f(x) = 1 − x if x ≤ 1 x + 1 if 1 < x < 3 x2 − 5 if x ≥ 3

OpenStudy (anonymous):

evaluate the limit as x approaches 1 from the left and right. same thing with 3. if limits are different approaching from left and right, function is discontinuous

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1-} 1-x =0\] \[\lim_{x \rightarrow 1+} x+1 =2\] discontinuous at x=1, repeat process for x=3

OpenStudy (anonymous):

Its 1

OpenStudy (anonymous):

What about? h(x) = 3 − 3x if x < 1 1 if x = 1 x + 3 if x > 1

OpenStudy (anonymous):

its 1

OpenStudy (anonymous):

f(x) = 1/x if x < 0 x if 0 ≤ x ≤ 4 4x − 3 if x > 4